OpenStudy (anonymous):

Er... Someone please explain this? Ine^1? What is the In? How do I solve an equation like this one?

OpenStudy (michele_laino):

\(\ln\) stands for \(natural\) \(logarithm\) or logarithm whose base is the irrational number \(e=2.718281828459045...\) so, we can write: \[\Large \ln {e^1} = k\] where \(k\) is such that: \[\Large {e^k} = {e^1}\] what is \(k\)?

OpenStudy (michele_laino):

more precisely: the subsequent formula: \[\huge \ln {e^x} = y\] where \(x,y\) are two appropriate real numbers, is, according to the definition of logarithm, equivalent to this one: \[\huge {e^y} = {e^x}\]

OpenStudy (michele_laino):

the last equation, holds, if and only if: \[\huge y = x\]

OpenStudy (anonymous):

OH I feel so dumb... I get it, sorry. :D That was actually really simple.